OpenStudy (anonymous):

Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept: form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

OpenStudy (anonymous):

point slope form:: point is (x1,y1) and slope is m,,then the equation of the line is: y-y1=m(x-x1)

OpenStudy (anonymous):

so in this case the equation is:: y-6=-4(x+1) y-6=-4x-4 y+4x=-4+6 4x+y=2

OpenStudy (anonymous):

Thanks Thanks Thanks!!!!

OpenStudy (anonymous):

slope-intercept form:: y=mx+b This is called the slope-intercept form because "m" is the slope and "b" gives the y-intercept and (x,y) the required point.....